# Analytic Computation of Dilaton Black Hole Quasinormal Modes via Seiberg-Witten Theory

### v1

@@ I Introduction
Previous analyses of EMd or related dilatonic black holes have mostly employed conventional techniquesMalik:2024sxv ; Konoplya:2022zav , whereas our work establishes a direct link to the quantum Seiberg-Witten (SW) framework. In $\mathcal{N}=2$ supersymmetric Yang-Mills theory, the SW construction provides exact results for strongly coupled dynamics, with the quantum SW curve encoding non-perturbative physics Seiberg:1994rs ; Seiberg:1994aj . For the case with $N_{f}=3$ hypermultiplets, the quantum SW curve can be cast into the form of the confluent Heun equation. Similar structures arise in black hole perturbation theory: the radial equation for a massive scalar field in an EMd background can also be expressed in confluent Heun form. This parallelism suggests a novel correspondence between black hole perturbations and the quantum SW curve, opening new opportunities to connect gravitational dynamics with gauge-theoretic methods.Bonelli:2021uvf ; Aminov:2020yma ; Bianchi:2021mft ; Ge:2025yqk ; Ge:2024jdx ; Silva:2025khf
In this work, we establish a concrete correspondence between the quantum Seiberg-Witten (SW) curve with $N_{f}=3$ and the confluent Heun equation, and further demonstrate its connection to the perturbation equations of Einstein-Maxwell-dilaton (EMd) black holes. By employing the normal form of the confluent Heun equation as a bridge, we show that the QNM problem in the EMd background can be reformulated in gauge-theoretic terms. We also analyze the boundary conditions relevant to both the SW curve and black hole perturbations, and compute the QNM spectrum using the WKB and continued fraction methods. These results not only deepen the understanding of QNMs in dilatonic black hole spacetimes but also reveal a new aspect of the gauge/gravity correspondence.
The paper is organized as follows. In Section II, we derive the field equations of Einstein-Maxwell-dilaton black holes. In Section III, we construct the correspondence between the quantum SW curve and the perturbation equations of EMd black holes via the confluent Heun form, and derive the QNM spectrum by imposing the QNM boundary conditions together with the SW quantization condition. Section IV presents the numerical QNM spectra of EMd black holes obtained from the quantum SW approach, benchmarked against the results from the WKB and continued fraction methods. Finally, Section V summarizes our findings and confirms the accuracy and effectiveness of the SW framework in capturing the quasinormal dynamics of EMd black holes.
@@ II Basic equations
In this work, we focus on the Einstein-Maxwell-dilaton (EMd) gravity, whose dynamics are governed by the action
@@
For radial equation Eq. (8), in order to establish a correspondence with the quantum Seiberg-Witten curve, we need to transform it into the normal form of the confluent Heun equation. We perform the following coordinate transformation and wavefunction rescaling
@@
III Dictionary and connection formula
In this section, we employ the normal form of the confluent Heun equation as a bridge to establish the correspondence between the quantum Seiberg-Witten curve and the equations of the Einstein-Maxwell-dilaton system. In addition, we discuss the associated boundary conditions.
III.1 Dictionay between Seiberg-Witten curve and Dilaton Black Hole
Seiberg-Witten theory offers an exact solution to four-dimensional $\mathcal{N}=2$ supersymmetric Yang-Mills theory, describing the moduli space of vacua and the low-energy dynamics through geometric methods, with significant implications for strong coupling and dualities. In this work, we mainly focus on the SU(2) $\mathcal{N}=2$ supersymmetric Yang-Mills theory with $N_{f}=3$ hypermultiplets (hypermatter fields). The corresponding quantum Seiberg-Witten curve can be expressed in the form of a confluent Heun equationAminov:2020yma :
@@ III.1 Dictionay between Seiberg-Witten curve and Dilaton Black Hole
On the torus, the integration contours can be chosen along the A-cycles and B-cycles. The corresponding quantization condition for the Seiberg-Witten spectrum reads
@@
where $\mathcal{F}^{(N_{f})}$ denotes the NS free energy of the SU(2) theory.
@@ III.2 Connection formula
we obtain the quasinormal modes. The Nekrasov-Shatashvili (NS) free energy can be obtained by instanton counting techniques (see Appendix A). The derivation of the connection formula and the quantization condition is presented in Appendix B.
@@ IV Numerical results
We computed the quasinormal modes (QNMs) of the dilaton black hole under various parameter settings using the quantum Seiberg-Witten (SW) quantization condition. In this work, the calculations were carried out up to $\Lambda_{3}^{4}$. To verify the validity and accuracy of our results, we further employed two well-established semi-analytical methods: the WKB approximation Schutz:1985km ; Iyer:1986np ; Konoplya:2003ii ; Matyjasek:2017psv ; Konoplya:2019hlu and the continued fraction method (CFM) Leaver:1985ax ; Konoplya:2011qq , both widely used in the study of black hole quasinormal modes. The detailed computation procedure for the continued fraction method is presented in Appendix C.
@@
Table 1 and Table 2 list the quasinormal frequencies for $l=1$ and $l=2$, respectively, obtained using the quantum SW, WKB, and continued fraction methods for different values of the scalar field mass $M\mu=0.01,0.1,0.2$ and black hole charge $Q=0.01,0.2,0.4$. The three approaches exhibit excellent consistency across all parameter sets, confirming the reliability of the quantum SW framework in reproducing the QNM spectra of dilaton black holes.
@@ V Summary and discussion
In this work, we studied the quasinormal modes of dilaton black holes in Einstein-Maxwell-dilaton gravity by mapping the scalar field equation to the quantum Seiberg-Witten curve of SU(2) $\mathcal{N}=2$ gauge theory with $N_{f}=3$. This correspondence allowed us to derive the QNM spectrum from the quantum SW quantization condition, yielding results in excellent agreement with the WKB method and the CFM method, with discrepancies typically below $10^{-3}$. The frequencies exhibit consistent trends: larger charge $Q$ or scalar mass $M\mu$ enhance the oscillation frequency, while increasing $M\mu$ reduces the damping rate, and higher angular momentum $l$ leads to longer-lived modes.
Compared with traditional approaches such as the WKB or continued-fraction methods, the quantum Seiberg-Witten framework provides a geometric and analytic formulation that interprets quasinormal mode spectra as quantized periods and systematically incorporates nonperturbative and instanton corrections, revealing the global and field-theoretic structure of the spectrum beyond purely numerical computations.
@@
Looking ahead, it would be interesting to extend the present analysis in several directions. Since the $N_{f}=3$ quantum Seiberg-Witten curve can be written as a confluent Heun equation, one natural step is to explore the cases $N_{f}=0,1,2$, whose curves can be expressed in terms of special forms of the biconfluent Heun equation. Some of these may be connected to the perturbation equations of black branes such as D3Bianchi:2021xpr ; Imaizumi:2022qbi , M5Imaizumi:2022dgj , and M2 branes, though their boundary conditions are typically more involved and deserve further study. In addition, we plan to investigate the quasinormal modes of dilaton black holes under different types of perturbations, such as Proca and Dirac fields, in order to better understand the universality and robustness of the gauge/gravity correspondence revealed in this work.
@@ Appendix A Instanton counting
In this appendix we provide a brief account of the computation of the Nekrasov-Shatashvili (NS) free energy using instanton counting techniques Nekrasov:2002qd ; Nekrasov:2009rc . This approach, originally developed in the context of four-dimensional $\mathcal{N}=2$ supersymmetric gauge theories, allows for a systematic evaluation of the instanton contributions to the prepotential and its $\Omega$-deformed extensions.
The Nekrasov-Shatashvili free energy of the SU(2) theory is composed of the classical, one-loop, and instanton contributions. Its derivative with respect to the quantum mirror map $a$ admits an explicit closed-form expression
@@ Appendix B Connection formula and quantization condition
To compute the quasinormal modes, we impose the condition that the ingoing wave at infinity vanishes. This requirement leads to $C_{1}(\Lambda,a,\mathbf{m})=0$, yielding
@@
In the Nekrasov-Shatashvili (NS) limit, we have
@@ Appendix C Continued fraction method
In this section, we present the detailed procedure for computing the quasinormal modes (QNMs) in the background of a dilaton black hole using the continued fraction method. We first introduce the coordinate transformation

### v2

@@ I Introduction
Previous analyses of EMd or related dilatonic black holes have mostly employed conventional techniquesMalik:2024sxv ; Konoplya:2022zav ; Ferrari:2000ep ; Brito:2018hjh ; Pope:2024ncb , whereas our work establishes a direct link to the quantum Seiberg-Witten (SW) framework. In $\mathcal{N}=2$ supersymmetric Yang-Mills theory, the SW construction provides exact results for strongly coupled dynamics, with the quantum SW curve encoding non-perturbative physics Seiberg:1994rs ; Seiberg:1994aj . For the case with $N_{f}=3$ hypermultiplets, the quantum SW curve can be cast into the form of the confluent Heun equation. Similar structures arise in black hole perturbation theory: the radial equation for a massive scalar field in an EMd background can also be expressed in confluent Heun form. This parallelism suggests a novel correspondence between black hole perturbations and the quantum SW curve, opening new opportunities to connect gravitational dynamics with gauge-theoretic methods.Bonelli:2021uvf ; Aminov:2020yma ; Bianchi:2021mft ; Ge:2025yqk ; Ge:2024jdx ; Silva:2025khf ; Lei:2023mqx
In this work, we establish a concrete correspondence between the quantum SW curve with $N_{f}=3$ and the confluent Heun equation, and further demonstrate its connection to the perturbation equations of EMd black holes. By employing the normal form of the confluent Heun equation as a bridge, we show that the QNM problem in the EMd background can be reformulated in gauge-theoretic terms. We also analyze the boundary conditions relevant to both the SW curve and black hole perturbations, and compute the QNM spectrum using the WKB and continued fraction methods. These results not only deepen the understanding of QNMs in dilatonic black hole spacetimes but also reveal a new aspect of the gauge/gravity correspondence.
The paper is organized as follows. In Section II, we derive the field equations of EMd black holes. In Section III, we construct the correspondence between the quantum SW curve and the perturbation equations of EMd black holes via the confluent Heun form, and derive the QNM spectrum by imposing the QNM boundary conditions together with the SW quantization condition. Section IV presents the numerical QNM spectra of EMd black holes obtained from the quantum SW approach, benchmarked against the results from the WKB and continued fraction methods. Finally, Section V summarizes our findings and confirms the accuracy and effectiveness of the SW framework in capturing the quasinormal dynamics of EMd black holes.
@@ II Basic equations
In this work, we focus on the EMd gravity, whose dynamics are governed by the action
@@
For radial equation Eq. (8), in order to establish a correspondence with the quantum SW curve, we need to transform it into the normal form of the confluent Heun equation. We perform the following coordinate transformation and wavefunction rescaling
@@
III Correspondence and connection formula
In this section, we employ the normal form of the confluent Heun equation as a bridge to establish the correspondence between the quantum SW curve and the equations of the EMd system. In addition, we discuss the associated boundary conditions.
III.1 Correspondence between Seiberg-Witten curve and Dilaton Black Hole
The SW theory offers an exact solution to four-dimensional $\mathcal{N}=2$ supersymmetric Yang-Mills theory, describing the moduli space of vacua and the low-energy dynamics through geometric methods, with significant implications for strong coupling and dualities. In this work, we mainly focus on the SU(2) $\mathcal{N}=2$ supersymmetric Yang-Mills theory with $N_{f}=3$ hypermultiplets (hypermatter fields). The corresponding quantum SW curve can be expressed in the form of a confluent Heun equationAminov:2020yma :
@@ III.1 Correspondence between Seiberg-Witten curve and Dilaton Black Hole
On the torus, the integration contours can be chosen along the A-cycles and B-cycles. The corresponding quantization condition for the SW spectrum reads
@@
where $\mathcal{F}^{(N_{f})}$ denotes the Nekrasov-Shatashvili (NS) free energy of the SU(2) theory.
@@ III.2 Connection formula
we obtain the QNMs. The NS free energy can be obtained by instanton counting techniques (see Appendix A). The derivation of the connection formula and the quantization condition is presented in Appendix B.
@@ IV Numerical results
We computed the QNMs of the dilaton black hole under various parameter settings using the quantum SW quantization condition. In this work, the calculations were carried out up to $\Lambda_{3}^{4}$. To verify the validity and accuracy of our results, we further employed two well-established semi-analytical methods: the WKB approximation Schutz:1985km ; Iyer:1986np ; Konoplya:2003ii ; Matyjasek:2017psv ; Konoplya:2019hlu and the continued fraction method (CFM) Leaver:1985ax ; Konoplya:2011qq , both widely used in the study of black hole QNMs. The detailed computation procedure for the CFM is presented in Appendix C.
@@
Table 1 and Table 2 list the quasinormal frequencies for $l=1$ and $l=2$, respectively, obtained using the quantum SW, WKBWKBpackage , and continued fraction methods for different values of the scalar field mass $M\mu=0.01,0.1,0.2$ and black hole charge $Q=0.01,0.2,0.4$. The three approaches exhibit excellent consistency across all parameter sets, confirming the reliability of the quantum SW framework in reproducing the QNM spectra of dilaton black holes.
@@ V Summary and discussion
In this work, we studied the QNMs of dilaton black holes in EMd gravity by mapping the scalar field equation to the quantum SW curve of SU(2) $\mathcal{N}=2$ gauge theory with $N_{f}=3$. This correspondence allowed us to derive the QNM spectrum from the quantum SW quantization condition, yielding results in excellent agreement with the WKB method and the CFM method, with discrepancies typically below $10^{-3}$. The frequencies exhibit consistent trends: larger charge $Q$ or scalar mass $M\mu$ enhance the oscillation frequency, while increasing $M\mu$ reduces the damping rate, and higher angular momentum $l$ leads to longer-lived modes.
Compared with traditional approaches such as the WKB or continued-fraction methods, the quantum SW framework provides a geometric and analytic formulation that interprets QNM spectra as quantized periods and systematically incorporates nonperturbative and instanton corrections, revealing the global and field-theoretic structure of the spectrum beyond purely numerical computations.
@@
Looking ahead, it would be interesting to extend the present analysis in several directions. Since the $N_{f}=3$ quantum SW curve can be written as a confluent Heun equation, one natural step is to explore the cases $N_{f}=0,1,2$, whose curves can be expressed in terms of special forms of the biconfluent Heun equation. Some of these may be connected to the perturbation equations of black branes such as D3Bianchi:2021xpr ; Imaizumi:2022qbi , M5Imaizumi:2022dgj , and M2 branes, though their boundary conditions are typically more involved and deserve further study. In addition, we plan to investigate the QNMs of dilaton black holes under different types of perturbations, such as Proca and Dirac fields, in order to better understand the universality and robustness of the gauge/gravity correspondence revealed in this work.
@@ Appendix A Instanton counting
In this appendix we provide a brief account of the computation of the NS free energy using instanton counting techniques Nekrasov:2002qd ; Nekrasov:2009rc . This approach, originally developed in the context of four-dimensional $\mathcal{N}=2$ supersymmetric gauge theories, allows for a systematic evaluation of the instanton contributions to the prepotential and its $\Omega$-deformed extensions.
The NS free energy of the SU(2) theory is composed of the classical, one-loop, and instanton contributions. Its derivative with respect to the quantum mirror map $a$ admits an explicit closed-form expression
@@ Appendix B Connection formula and quantization condition
To compute the QNMs, we impose the condition that the ingoing wave at infinity vanishes. This requirement leads to $C_{1}(\Lambda,a,\mathbf{m})=0$, yielding
@@
In the NS limit, we have
@@ Appendix C Continued fraction method
In this section, we present the detailed procedure for computing the QNMs in the background of a dilaton black hole using the CFM. We first introduce the coordinate transformation